i have another solve for the unknown
8/52 = m/39
again i know this is basic math but i am lost any help please thank you

Found 2 solutions by blueskyy, Annabelle1:
Answer by blueskyy(3)



8/52 = M/39
SO, (8 * 39)/ 52= M
SO, 6 = M

Answer by Annabelle1(69)

8/52 = m/39
this one is similar to the last. we wamt m on its own so we need to cancel out the /39. we do this my multiplying both sides of the equation by 39
(8/52)*39=(m/39)*39
m/39*39 is m so we have
(8/52)*39=m
6=m
m=6